# Galactic Archaeology with Gaia

### v1

@@ 5.3 Census of accreted substructures
What is the origin of the multitude of halo substructures discovered by Gaia in the form of IOM clumps (see Fig. 23) or as coherent narrow streams (see Fig. 24)? More specifically, how many of these sub-structures are tidal remnants of destroyed star clusters and how many come from dwarf galaxies? A direct estimate of the total number of star cluster streams can be found in Baumgardt et al. (2019). These authors show that the initial mass function of the survived and currently observable Milky Way globular clusters is a very strong function of their Galacto-centric distance (see Fig. 25). Assuming that the cluster initial mass function should not depend strongly on the current location in the Galaxy, the total number of destroyed GCs can be estimated. Such an estimate depends on the postulated shape of the initial GC mass function. Baumgardt et al. (2019) consider two possibilities: the log-normal mass function, resembling the overall current distribution of the GC masses, and the power-law mass function, usually preferred as a model of the cluster mass distribution at birth. In the former case, roughly as many GCs are destroyed as there are currently detected ($N\approx 200$), but in the latter, the total initial number is much larger ($N\approx 10,000$). Irrespective of the initial mass function parameterisation, the bulk of the destroyed GCs are located within the Solar radius, with a large fraction within $\approx 5$ kpc from the Galactic centre, in the region matching in extent the in-situ Aurora component of the Milky Way.
@@
In the aftermath of the Gaia data boon it is worth considering if we now have a good handle on all (or at least most) of the relevant accreted structures in the halo. The short answer is almost certainly not, but we have likely uncovered the most significant progenitors by mass. This deduction is clearly laid out by Naidu et al. (2020), who use Gaia and spectroscopic data from the H3 survey to estimate the relative fraction of structures in the halo as a function of distance (see Fig. 26.) The authors argue that the GS/E dominates in the inner halo ($r\lesssim 20$ kpc), and the majority of stars in the outer halo belong to Sgr. Of the other structures considered in Naidu et al. (2020), none contribute more than 5% at any distance!
@@
Figure 26: Relative fraction of structures in the Milky Way as a function of Galactocentric distance [reproduced from Naidu et al. (2020)]. These fractions were estimated using Gaia and H3 survey data. Of the accreted structures, GS/E dominates in the inner halo, and Sgr dominates the outer halo. No other substructure contributes more than $\sim 5\%$ at any distance.
@@
Theoretical predictions for the expected number of destroyed dwarf galaxies in Milky Way-mass haloes are relatively scarce, but notably Fattahi et al. (2020) show that the number of destroyed dwarfs generally exceeds the number of surviving dwarfs at all mass scales (see Figure 1 in Fattahi et al. 2020). Moreover, as expected, there can be significant halo-to-halo scatter due to varying mass accretion histories (see also Deason et al. 2023). Attempts to quantify the mass spectrum of accreted dwarfs in the Milky Way data are few and far between, but notably dynamical grouping (e.g. Callingham et al., 2022) and/or the use of metallicity distribution functions or chemical planes (e.g. Cunningham et al., 2022; Deason et al., 2023) are likely the best tools we have moving forward. Deason et al. (2023) use GMM modeling of the stellar halo MDF to estimate the luminosity function of destroyed dwarfs in the Milky Way (see Fig. 27). They make use of the stellar mass-metallicity relation and assume Gaussian MDF distributions for individual progenitors, thus assuming the overall stellar halo MDF is a mixture of MDFs from smaller galaxies. Applying this method to a hodge podge of spectroscopic data in the Milky Way, complemented by Gaia DR3 astrometry, indicates that that the Milky Way stellar halo has $N\sim 1-3$ massive progenitors (with $L>10^{8}L_{\odot}$) within 10 kpc, and likely several hundred progenitors in total.
@@
Figure 28: Left panel: [Fe/H] vs. stellar mass at infall for destroyed (red points) and surviving (blue points) dwarfs in the Auriga simulations. The average relations at fixed stellar mass are shown with the solid lines. For comparison, the dashed line shows the average relation for $z=0$ satellites. [Adapted from Fattahi et al. (2020)]. Right panel: [Fe/H] vs. stellar mass for surviving (gray) and disrupted (purple) dwarfs in the Milky Way. The gray line is the $z=0$ stellar mass metallicity from Kirby et al. (2013). The purple line is offset from this relation by 0.3 dex, which is the median offset of the disrupted dwarfs.[Adapted from Naidu et al. (2022)].
An important consideration when attempting to model the MDF of halo stars is the evolution of the MZR with time (see also earlier discussion about how this relates to the GS/E properties in Section 2). The redshift $z=0$ MZR for surviving satellites in the Milky Way is well-documented (Kirby et al., 2013), however, can we simply assume that this relation holds at all redshifts? Fig. 27 shows how modeling of the stellar halo MDF is significantly impacted by the choice of MZR, and this is thus an important caveat in any such work. At the dwarf mass-scales, both observations and simulations seem to suggest that there is evolution of the MZR over time (see Fig. 28). Fattahi et al. (2020) and Naidu et al. (2022) show that destroyed dwarfs are metal-poorer (by $\sim 0.3$ dex) than their surviving dwarf counterparts at fixed stellar mass. This difference is likely owing to the different star formation timescales of the two populations. Destroyed dwarfs are typically accreted earlier than surviving dwarfs, and thus to get to the same stellar mass as their $z=0$ counterparts, destroyed dwarfs must have formed their stars on a shorter timescale, and thus have less time to build-up their metals. In support of this picture, Naidu et al. (2022) also show that the destroyed dwarfs are more $\alpha$-enhanced than surviving dwarfs at fixed stellar mass.
@@ 6 Conclusions and future outlook
The Gaia era has revolutionised the field of Galactic Archaeology. It has changed our view on how the Milky Way assembled, it has changed how we model the Galaxy, and it has opened our eyes to new research directions and explorations. However, the era is not over yet! In the coming years, the final Gaia data releases will become available and our Gaia all-sky map of the Galaxy will be complete. This, coupled with the complementary spectroscopic surveys such as DESI, WEAVE, 4MOST, SDSS-V, etc will provide the next phase in our Galactic exploration. Looking further ahead, pushing these missions to fainter magnitudes, and thus greater distances will be paramount (e.g. GaiaNIR).
In the wise words of Donald Lynden-Bell, one should always try to “follow the data”. That, given the quality and abundance of data at our fingertips, should be advice well heeded. However, the interpretation of this data requires detailed modeling and simulations that will need to keep pace with the data in order to achieve the ultimate goal: a multi-dimensional, star-by-star ‘replay’ of the Galaxy’s formation.
Below we list some of the most important and puzzling unanswered questions in Galactic Archaeology that future datasets and modeling efforts will be striving to address:
@@ Acknowledgments
AD is supported by a Royal Society University Research Fellowship. AD acknowledges support from the Leverhulme Trust and the Science and Technology Facilities Council (STFC) [grant numbers ST/P000541/1, ST/T000244/1]. VB acknowledges support from the Leverhulme Research Project Grant RPG-2021-205: ”The Faint Universe Made Visible with Machine Learning”.

### v2

@@ 5.3 Census of accreted substructures
What is the origin of the multitude of halo substructures discovered by Gaia in the form of IOM clumps (see Fig. 23) or as coherent narrow streams (see Fig. 24)? More specifically, how many of these sub-structures are tidal remnants of destroyed star clusters and how many come from dwarf galaxies? A direct estimate of the total number of star cluster streams can be found in Baumgardt et al. (2019). These authors show that the initial mass function of the survived and currently observable Milky Way globular clusters is a very strong function of their Galacto-centric distance (see Fig. 25). Assuming that the cluster initial mass function should not depend strongly on the current location in the Galaxy, the total number of destroyed GCs can be estimated. Such an estimate depends on the postulated shape of the initial GC mass function. Baumgardt et al. (2019) consider two possibilities: the log-normal mass function, resembling the overall current distribution of the GC masses, and the power-law mass function, usually preferred as a model of the cluster mass distribution at birth. In the former case, roughly as many GCs are destroyed as there are currently detected ($N\approx 200$), but in the latter, the total initial number is much larger ($N\approx 10,000$). Irrespective of the initial mass function parameterization, the bulk of the destroyed GCs are located within the Solar radius, with a large fraction within $\approx 5$ kpc from the Galactic centre, in the region matching in extent the in-situ Aurora component of the Milky Way.
Many of the stellar streams recently discovered with the STREAMFINDER algorithm (Ibata et al., 2021, and Fig 24) have small intrinsic widths (by design) and are therefore suspected to have originated from disrupted or disrupting globular clusters (see also Martin et al., 2022a). Of these, the C-19 system is particularly striking. It is kinematically cold (although see Yuan et al., 2022, for a follow-up study), and displays chemical anomalies characteristic of GCs, with a large spread and correlations amongst abundances of Mg, Na and Al (Martin et al., 2022b; Yuan et al., 2022). What makes it stand out compared to known GCs is its extreme metallicity: at [Fe/H]$\approx-3.4$, C-19 would be the most metal-poor GC known to date in the Milky Way, by far (Martin et al., 2022b). C-19 could plausibly be a harbinger of a previously undetected population of metal-poor GCs, indicating that the so-called GC metallicity floor is, at least in part, due to a selection bias. Moreover, as pointed out by Martin et al. (2022b), given the typical stellar metallicity distribution function, the fractional contribution of GCs to star-formation in the C-19’s host galaxy must be high. This is in agreement with the most recent constraints in the ancient Milky Way (Belokurov and Kravtsov, 2023) and in high-redshift galaxies (Mowla et al., 2024).
@@
In the aftermath of the Gaia data boon it is worth considering if we now have a good handle on all (or at least most) of the relevant accreted structures in the halo. The short answer is almost certainly not, but we have likely uncovered the most significant progenitors by mass. This deduction is clearly laid out by Naidu et al. (2020), who use Gaia and spectroscopic data from the H3 survey to estimate the relative fraction of structures in the halo as a function of distance (see Fig. 26.) The authors argue that the GS/E dominates in the inner halo ($r\lesssim 20$ kpc), and the majority of stars in the outer halo belong to Sgr. Of the other structures considered in Naidu et al. (2020), none contribute more than 5% beyond $r\gtrsim 5$ kpc!
@@
Figure 26: Relative fraction of structures in the Milky Way as a function of Galactocentric distance. These fractions were estimated using Gaia and H3 survey data. Of the accreted structures, GS/E dominates in the inner halo, and Sgr dominates the outer halo. No other substructure contributes more than $\sim 5\%$ beyond $r\gtrsim 5$ kpc. [Reproduced from Naidu et al. (2020)]
@@
Theoretical predictions for the expected number of destroyed dwarf galaxies in Milky Way-mass haloes are relatively scarce, but notably Fattahi et al. (2020) show that the number of destroyed dwarfs generally exceeds the number of surviving dwarfs at all mass scales (see Figure 1 in Fattahi et al. 2020). Moreover, as expected, there can be significant halo-to-halo scatter due to varying mass accretion histories (see also Deason et al. 2023). Attempts to quantify the mass spectrum of accreted dwarfs in the Milky Way data are few and far between, but notably dynamical grouping (e.g. Callingham et al., 2022) and/or the use of metallicity distribution functions or chemical planes (e.g. Cunningham et al., 2022; Deason et al., 2023) are likely the best tools we have moving forward. Deason et al. (2023) use GMM modelling of the stellar halo MDF to estimate the luminosity function of destroyed dwarfs in the Milky Way (see Fig. 27). They make use of the stellar mass-metallicity relation and assume Gaussian MDF distributions for individual progenitors, thus assuming the overall stellar halo MDF is a mixture of MDFs from smaller galaxies. Applying this method to a hodge podge of spectroscopic data in the Milky Way, complemented by Gaia DR3 astrometry, indicates that that the Milky Way stellar halo has $N\sim 1-3$ massive progenitors (with $L>10^{8}L_{\odot}$) within 10 kpc, and likely several hundred progenitors in total.
@@
Figure 28: Left panel: [Fe/H] vs. stellar mass at infall for destroyed (red points) and surviving (blue points) dwarfs in the Auriga simulations (Fattahi et al., 2020). The average relations at fixed stellar mass are shown with the solid lines. For comparison, the dashed line shows the average relation for $z=0$ satellites. Right panel: [Fe/H] vs. stellar mass for surviving (grey) and disrupted (purple) dwarfs in the Milky Way (Naidu et al., 2022). The grey line is the $z=0$ stellar mass metallicity from Kirby et al. (2013). The purple line is offset from this relation by 0.3 dex, which is the median offset of the disrupted dwarfs. [Adapted from Fattahi et al. (2020); Naidu et al. (2022)].
An important consideration when attempting to model the MDF of halo stars is the evolution of the MZR with time (see also earlier discussion about how this relates to the GS/E properties in Section 2). The redshift $z=0$ MZR for surviving satellites in the Milky Way is well-documented (Kirby et al., 2013), however, can we simply assume that this relation holds at all redshifts? Fig. 27 shows how modelling of the stellar halo MDF is significantly impacted by the choice of MZR, and this is thus an important caveat in any such work. At the dwarf mass-scales, both observations and simulations seem to suggest that there is evolution of the MZR over time (see Fig. 28). Fattahi et al. (2020) and Naidu et al. (2022) show that destroyed dwarfs are metal-poorer (by $\sim 0.3$ dex) than their surviving dwarf counterparts at fixed stellar mass. This difference is likely owing to the different star formation timescales of the two populations. Destroyed dwarfs are typically accreted earlier than surviving dwarfs, and thus to get to the same stellar mass as their $z=0$ counterparts, destroyed dwarfs must have formed their stars on a shorter timescale, and thus have less time to build-up their metals. In support of this picture, Naidu et al. (2022) also show that the destroyed dwarfs are more $\alpha$-enhanced than surviving dwarfs at fixed stellar mass. However, it is worth noting that the nature of some of the ‘destroyed dwarfs’ used in the Naidu et al. (2022) work are still under debate (e.g. some of these apparently distinct substructures may be part of the GS/E). Thus, while there is reason to believe there is indeed evolution of the MZR at the dwarf mass scale, there is, as yet, no robust observational measure of this evolving MZR.
@@ 6 Conclusions and future outlook
The Gaia era has revolutionized the field of Galactic Archaeology. It has changed our view on how the Milky Way assembled, it has changed how we model the Galaxy, and it has opened our eyes to new research directions and explorations. However, the era is not over yet! In the coming years, the final Gaia data releases will become available and our Gaia all-sky map of the Galaxy will be complete. This, coupled with the complementary spectroscopic surveys such as DESI, WEAVE, 4MOST, SDSS-V, etc will provide the next phase in our Galactic exploration. Looking further ahead, pushing these missions to fainter magnitudes, and thus greater distances will be paramount (e.g. GaiaNIR).
In the wise words of Donald Lynden-Bell, one should always try to “follow the data”. That, given the quality and abundance of data at our fingertips, should be advice well heeded. However, the interpretation of this data requires detailed modelling and simulations that will need to keep pace with the data in order to achieve the ultimate goal: a multi-dimensional, star-by-star ‘replay’ of the Galaxy’s formation.
Below we list some of the most important and puzzling unanswered questions in Galactic Archaeology that future datasets and modelling efforts will be striving to address:
@@ Acknowledgments
AD is supported by a Royal Society University Research Fellowship. AD acknowledges support from the Leverhulme Trust and the Science and Technology Facilities Council (STFC) [grant numbers ST/X001075/1, ST/T000244/1]. VB acknowledges support from the Leverhulme Research Project Grant RPG-2021-205: “The Faint Universe Made Visible with Machine Learning”.
We thank an anonymous referee for providing a constructive report.